Documenta Mathematica

Band 14, 2009

Jaya NN Iyer, Stefan Müller–Stach Chow–Künneth Decomposition

for Some Moduli Spaces  1–18

Dennis Gaitsgory and David Nadler Hecke Operators on Quasimaps

into Horospherical Varieties  19–46

Mark L. MacDonald Projective Homogeneous Varieties

Birational to Quadrics  47–66

Ajay C. Ramadoss On the Nonexistence of Certain Morphisms from Grassmannian to Grassmannian

in Characteristic 0  67–113

Matthias Huber Spectral Analysis of Relativistic Atoms –

Interaction with the Quantized Radiation Field  115–156

Sebastian Neumann

Rationally Connected Foliations on Surfaces  157–165

Daniel Lenz, Peter Stollmann, Ivan Veselić The Allegretto-Piepenbrink Theorem

for Strongly Local Dirichlet Forms  167–189

Jan Nekovář Erratum for “On the Parity of Ranks of Selmer Groups III”

cf. Documenta Math. 12 (2007), 243–274  191–194

Marianne Akian, Stéphane Gaubert, and Cormac Walsh

The Max-Plus Martin Boundary  195–240

Christopher Skinner A note on the p-adic Galois representations

attached to Hilbert modular forms  241–258

Jonas Bergström Equivariant Counts of Points of the

Moduli Spaces of Pointed Hyperelliptic Curves  259–296

Matthias Huber Spectral Analysis of Relativistic Atoms –

Dirac Operators with Singular Potentials  297–338

J. M. Douglass and G. Röhrle Homology of the Steinberg Variety

and Weyl Group Coinvariants  339–357

David Gepner and Victor Snaith On the Motivic Spectra Representing

Algebraic Cobordism and Algebraic K-Theory  359–396

Yichao Tian

p-Adic Monodromy of the Universal

Deformation of a HW-Cyclic Barsotti-Tate Group  397–440

Mark Kisin and Wei Ren

Galois Representations and Lubin-Tate Groups  441–461

Rupert L. Frank, Heinz Siedentop, and Simone Warzel The Energy of Heavy Atoms According to Brown

and Ravenhall: The Scott Correction  463–516

Moshe Jarden and Florian Pop

Function Fields of One Variable over PAC Fields  517–523

Balmer and Calms Geometric Description

of the Connecting Homomorphism for Witt Groups  525–550

Niko Naumann, Markus Spitzweck, Paul Arne Østvær

Motivic Landweber Exactness  551–593

Gelu Popescu Hyperbolic Geometry

on Noncommutative Balls  595–651

Benedictus Margaux Vanishing of Hochschild Cohomology for Affine Group Schemes and Rigidity of Homomorphisms

between Algebraic Groups  653–672

Christian Miebach, Karl Oeljeklaus

On Proper ℝ-Actions on Hyperbolic Stein Surfaces  673–689

Finnur Lárusson

Affine Simplices in Oka Manifolds  691–697

Christian Schlichtkrull

Thom Spectra that Are Symmetric Spectra  699–748

Marc A. Nieper-Wißkirchen Twisted Cohomology of the

Hilbert Schemes of Points on Surfaces  749–770

Toby Gee The Sato-Tate Conjecture

for Modular Forms of Weight 3  771–800

Orr Shalit and Baruch Solel

Subproduct Systems  801–868